# Learned radio interferometric imaging for varying visibility coverage Learned radio interferometric imaging for varying visibility coverage

### v1

@@ Abstract
With the next generation of interferometric telescopes, such as the Square Kilometre Array (SKA), the need for highly computationally efficient reconstruction techniques is particularly acute. The challenge in designing learned, data-driven reconstruction techniques for radio interferometry is that they need to be agnostic to the varying visibility coverages of the telescope, since these are different for each observation. Because of this, learned post-processing or learned unrolled iterative reconstruction methods must typically be retrained for each specific observation, amounting to a large computational overhead. In this work we develop learned post-processing and unrolled iterative methods for varying visibility coverages, proposing training strategies to make these methods agnostic to variations in visibility coverage with minimal to no fine-tuning. Learned post-processing techniques are heavily dependent on the prior information encoded in training data and generalise poorly to other visibility coverages. In contrast, unrolled iterative methods, which include the telescope measurement operator inside the network, achieve state-of-the-art reconstruction quality and computation time, generalising well to other coverages and require little to no fine-tuning. Furthermore, they generalise well to realistic radio observations and are able to reconstruct the high dynamic range of these images.
@@ 1 Introduction
To study radio sources in the Universe at high-resolution there is a need for radio telescopes with a very large aperture that cannot be achieved with a single dish. Instead, radio interferometry forms a telescope using multiple dishes or antennae spaced out over long distances to synthesise an aperture large enough to image the Universe in enough detail to uncover its mysteries. By combining the measurements from two of these antennae, which is referred to as a baseline, a so-called visibility can be measured, corresponding to one Fourier component of the sky. The inversion of this measurement process is an ill-posed problem as the radio interferometric telescope samples the Fourier domain non-uniformly, with a finite number of samples, and are contaminated by measurement noise. By projecting the acquired data to the image domain one recovers what is called a dirty image, a reconstruction of the target convolved with the point-spread function (PSF) of the telescope. Typically, various algorithms are used to reconstruct the image from acquired visibilities to yield a more accurate clean image of the true radio sky (e.g. CLEAN, Högbom, 1974).
@@
Recently, learned approaches have gained in popularity, using training data sets as prior information for reconstructions. One particular approach of these learned methods implements learned regularisation, replacing the conventional regularisers in optimisation methods with learned priors. The most popular and promising method is to use so-called Plug-and-Play (PnP, Venkatakrishnan et al., 2013) denoisers that can be trained on different data domains, foregoing the need for large domain-specific data sets. Since these methods retain the decoupling of the prior and measurement operator they are robust to changes in the measurement operator and measurement noise level. In particular, PnP priors have shown promising results in radio imaging (Terris et al., 2023; Liaudat et al., 2023; Terris et al., 2022; Dabbech et al., 2022; Wilber et al., 2023). Furthermore, recent work has demonstrated how PnP methods can be extended to provide efficient uncertainty quantification (Liaudat et al., 2023). However, these methods require iterative reconstruction and so remain relatively computationally demanding.
@@
Learned unrolled iterative methods typically closely resemble iterative optimisation techniques, while reducing the number of iterations with the measurement operator, have shown promise in both reconstruction performance and in robustness to changes in the measurement operator. These methods have also produced promising results in astronomical imaging (e.g. Mars et al., 2023) and strike a balance between speed and quality of reconstruction. Inclusion of the measurement operator in the network architecture allows the network to retain some robustness to changes and uncertainty in the measurement operator (Boink et al., 2019, 2020; Maier et al., 2019), however results are conditioned on the measurement operator used during training.
@@ 2.2 Inverse problem
which provides a trade off between the data fidelity term $\mathcal{L}(\boldsymbol{\Phi}\boldsymbol{x},\boldsymbol{y})$ and the regularisation term $\mathcal{S}(\boldsymbol{x})$, which encodes prior information of the images and is weighted by the regularisation parameter $\lambda$. In radio interferometry, a typical data fidelity term is the squared $\ell_{2}-\text{norm}$, i.e. $\mathcal{L}(\boldsymbol{\Phi}\boldsymbol{x},\boldsymbol{y})=\|\boldsymbol{\Phi}\boldsymbol{x}-\boldsymbol{y}\|_{\ell_{2}}^{2}$, and a widely used regularisation term is an $\ell_{1}$-norm on an overcomplete dictionary of wavelet bases, e.g. consisting of a Dirac basis and Debauchies wavelets bases (Daubechies, 1992), as used in the sparsity averaging reweighted analysis (SARA, Carrillo et al., 2012) algorithm. Alternatively, this problem can be formulated in the constrained setting
$\boldsymbol{x}^{\star}=\underset{\boldsymbol{x}\in\mathbb{R}^{N}}{\text{arg min}}\quad\mathcal{S}(\boldsymbol{x})\quad\text{s.t.}\quad\|\boldsymbol{\Phi}\boldsymbol{x}-\boldsymbol{y}\|_{\ell_{2}}^{2}\leq\epsilon.$ (8)
The optimisation problems in Equations 7 and 8 are typically solved using proximal optimisation techniques, such as the fast iterative shrinkage-thresholding algorithm (FISTA, Beck & Teboulle, 2009), the alternating direction method of multipliers (ADMM, Boyd, 2010), or the primal-dual hybrid gradient method (PDHG, Chambolle et al., 2018) and a variety of these methods have been applied to radio interferometry (e.g. Pratley et al., 2018, 2019a).
Learned reconstruction approaches aim to replace some parts of these traditional optimisation techniques with neural networks in order to incorporate a data-driven prior and/or to speed up the reconstruction process. Learned regularisation approaches use neural networks to capture the prior information in the variational problem. In particular, PnP methods replace the proximal problem formulation in proximal splitting algorithms with a denoiser (Venkatakrishnan et al., 2013; Ryu et al., 2019). These denoisers are typically trained on data sets outside of the target domain, e.g. natural images, and encode more general prior information about the denoising of images. The iterative updates based on both these more general priors and the data fidelity term in the optimisation problem allow these methods to generalise well to unseen data. While these methods are able to achieve high reconstruction quality they are computationally expensive as they require many iterations and therefore evaluations of the measurement operator to converge to a solution.
Instead, learned post-processing methods (e.g. Chen et al., 2017; Jin et al., 2017; Yi & Babyn, 2018) aim to drastically speed up the reconstruction process by training a neural network to post-process an initial reconstruction. In radio interferometry, this initial reconstruction is typically the dirty image obtained by the pseudo-inverse of the measurements. Figure 1 shows the typical pipeline of a learned post-processing method in radio interferometry. Because the network is decoupled from the measurement process, these methods can be trained using an image-to-image network making them much faster to train and evaluate. However, these methods are heavily reliant on the quality of the initial reconstruction and dependant on the artefact distribution of the uv-coverage(s) used for the training data.
@@
Learned unrolled iterative methods combine the data-driven priors with model based updates in order to provide a balance between the speed of learned post-processing methods and the quality of learned regularisation methods. They typically mimic proximal optimisation algorithms by unrolling a small number of iterations of the iterative process and replacing some of the proximal operators with neural networks (Gregor & LeCun, 2010). Figure 2 shows a diagram summarising these methods. While a large portion of these methods closely follow the traditional optimisation methods (e.g. Adler & Öktem, 2018), some methods have been proposed that incorporate the measurement process more directly into the network. For instance, Hauptmann et al. (2020) proposed a multi-scale learned iterative method that includes the measurement information at different resolution scales in the multi-scale U-Net architecture (Ronneberger et al., 2015). Similarly there are methods that see the inclusion of the gradient of a data fidelity term added into this U-Net structure at several resolution scales (Trent, 2020; Mars et al., 2023). The benefit in these methods is the reduced computation and memory overhead as at the lower scales of the network the measurement operator is evaluated at a lower resolution and therefore more efficiently.
@@ 3.2 Learned unrolled iterative method
The second approach we consider is a learned unrolled iterative method that uses the measurement operator and measurements of the observation in the neural network to improve the reconstruction quality. We extend the U-Net architecture by introducing measurement information at each scale of the network. To achieve this we use so-called sub-scale operators $\boldsymbol{\Phi}_{i}:\mathbb{R}^{N}_{i}\rightarrow\mathbb{C}^{M}_{i}$ that map from a reduced image space $\boldsymbol{x}_{i}\in\mathbb{R}^{N_{i}}$ to low-pass filtered visibilities $\boldsymbol{y}_{i}\in\mathbb{C}^{M_{i}}$. The use of sub-scale operators is discussed in more detail in Mars et al. (2023). These sub-scale operators are used to include the gradient of the data fidelity term $\mathcal{L}(\boldsymbol{\Phi}_{i}\boldsymbol{x}_{i},\boldsymbol{y}_{i})=\frac{1}{2}\|\boldsymbol{\Phi}_{i}\boldsymbol{x}_{i}-\boldsymbol{y}_{i}\|_{\ell_{2}}^{2}$ in the multi-scale U-Net:
@@ 4.2 Reconstruction quality
4.3 High dynamic range reconstruction quality
While the proposed methods are trained on images of simulated galaxies, the resulting networks should also be tested on more realistic radio observations. Therefore, to test the generalisation power of the different training strategies, we evaluate the performance of the models on an out-of-sample image of 30 Doradus. The dynamic range of this image is significantly larger than the simulated images and we therefore also evaluate the reconstructions on a logarithmic scale.
@@ 4.3 High dynamic range reconstruction quality
Figure 8: Reconstructions using the trained variants of the U-Net on simulated measurements, using the true uv-coverage drawn from a Gaussian random distribution, of an out-of-sample image of 30 Doradus, displayed on a linear scale (top) and a logarithmic scale (bottom). The strategies that are not fine-tuned on data from the true coverage perform significantly worse in reconstructing this out-of-sample image. Using transfer learning from the single coverage strategy brings the reconstructive SNR closer to fully retraining the model on the true coverage. All U-Net reconstructions struggle to accurately reconstruct the high dynamic range of this image as indicated by the logarithmically scaled images and the logSNR, with the single coverage strategy producing large artefacts.
When looking at the reconstructions from the same measurements using the GU-Net in Figure 9, we see that the models are able to produce high quality reconstructions with a high dynamic range. Similar to on the simulated images, the methods not fine-tuned on data from the true coverage performs slightly worse but this difference is easily mitigated by fine-tuning the models using transfer learning, yielding reconstructions that are of the same reconstruction quality as fully retraining the model. The explicit inclusion of the measurement operator in the reconstruction process allows the network to learn the prior information irrespective of the coverage used, which is why the GU-Net is able to generalise much better to out-of-distribution data as well as generalising to different uv-coverages. This generalisablity makes it more suitable for application to real radio observations, since it is able to correct for differences between the training data set and the actual observations.
Figure 9: Reconstructions using the trained variants of the GU-Net on simulated measurements, using the true uv-coverage drawn from a Gaussian random distribution, of an out-of-sample image of 30 Doradus, displayed on a linear scale (top) and a logarithmic scale (bottom). The GU-Net is able to reconstruct the high dynamic range of the image across the different strategies as indicated by both the high SNR and logSNR of the reconstructions. Training the model on a single coverage and then fine-tuning it using transfer learning provides reconstruction quality close to that of the model that is fully trained on simulated data from the true coverage.
@@ 4.4 Transfer learning to MeerKAT uv-coverage
Finally, to investigate whether these reconstruction networks generalise from the Gaussian visibility coverages to realistic uv-coverages, we use the same models to reconstruct from simulated measurements using a uv-coverage from the MeerKAT telescope with $241920$ visibilities as seen in Figure 10. In order to do this we fine-tune the GU-Net model trained on the single coverage by training it on a small data set created using the MeerKAT coverage. The model is trained for 50 epochs and the creation of the data took $\sim 2$ hours and the fine-tuning of the model took $\sim 5$ hours on an A100 GPU with 40GB VRAM. For this task we only consider the GU-Net architecture as the U-Net models do not generalise well to high dynamic range images as demonstrated in Section 4.3. The fine-tuned GU-Net is then used to reconstruct an out-of-sample image of 30 Doradus using the MeerKAT coverage. The results are shown in Figure 11.
@@
Figure 11: Reconstructions using a fine-tuned variant of the GU-Net on simulated measurements on an out-of-sample image of 30 Doradus using the MeerKAT uv-coverage as displayed in Figure 10. The ground truth, dirty image and GU-Net reconstruction are displayed on a linear scale (top) and a logarithmic scale (bottom). Using this fine-tuned model we achieve a high quality reconstruction that also reproduces the high dynamic range of the radio observation, using only minimal amount of re-training.
We see that the transfer learned model is capable of reconstructing images with high dynamic range, despite not being trained on any data with a similar dynamic range. Since the number of measurements in the MeerKAT coverage is larger than that of the Gaussian coverage (7.4$\sim$ 7.4$\times$ 7.4 × more measurements), the reconstruction time is longer ($194.6\pm 3.1$ms; 2.3$\sim$ 2.3$\times$ 2.3 × longer) than for the simulated measurements and the training takes longer.
@@ 5 Conclusions
First of all, we conclude that methods that do not include the measurement operator in the reconstruction network, such as learned post-processing methods based on a U-Net model, are not robust to changes in the uv-coverage and are highly reliant on learning features that are part of the training data. While these methods are the most computationally efficient reconstruction approaches, their reconstructive performance is significantly worse than methods that do include the measurement operator, even when fully (re)trained or fine-tuned to the true uv-coverage, due to overfitting. Furthermore, by looking at the performance of this post-processing approach on more realistic radio interferometric observations, we have shown that these methods struggle to capture the high dynamic range of these images across the different training strategies. This indicates that these methods are limited by the prior information in their training data, which is concerning in the context of real data where the true image is unknown.
@@
Furthermore, we have shown that when applying the GU-Net reconstruction method to measurements outside of the training domain, such as a more realistic radio observation with a high dynamic range, we retain this reconstruction quality, indicative of the approach being significantly less reliant on the prior information in the training data and able to generalise across different types of data.
@@
In order to train these models on distributions of realistic uv-coverages, some technical challenges need to be addressed regarding the methods of varying the coverage during training causing a large overhead in terms of memory and/or compilation time. During the training of our networks on random Gaussian coverages, we have addressed this by using a random subset of a larger coverage, which allows us to train on a distribution of coverages without the need to change the coverage during training, thus requiring only a minor increase in memory necessary. However, this is would be highly computationally demanding in the case of realistic uv-coverages. In order to address this, we propose to investigate the use of approximate measurement operators in the reconstruction network by using a convolution with the PSF to approximate the measurement operator as a way to reduce the computational overhead of changing the coverage during training, as has been investigated for proximal optimisation methods (Bester et al., 2021). While this approximation does not model wide-field effects when the baselines are non co-planar, this would allow us to train on a distribution of realistic uv-coverages without the need to change the coverage during training, which would be essential in order to train these networks. However, the difference in performance between using the full measurement operator and the approximate operator in learned methods should be investigated further to assess how much information is lost.
Another interesting avenue for future work would be to investigate not only the fine tuning on the true uv-coverage, but also the performance of transfer learning on real data. In Mars et al. (2023) we have shown that transfer learning can be used to improve the performance of our reconstruction methods trained on a general image data set, by using a small amount of domain specific data. This approach could bridge the gap between a larger data set of simulated data and a small data set of real data.
Finally, a more extensive investigation should be done with regards to the performance of these methods on even more realistic radio observations, including beam errors, calibration errors and other errors that might be present in the actual data, as well as including bright point sources in the data sets which are not present in the current training data set.

### v2

@@ Abstract
With the next generation of interferometric telescopes, such as the Square Kilometre Array (SKA), the need for highly computationally efficient reconstruction techniques is particularly acute. The challenge in designing learned, data-driven reconstruction techniques for radio interferometry is that they need to be agnostic to the varying visibility coverages of the telescope, since these are different for each observation. Because of this, learned post-processing or learned unrolled iterative reconstruction methods must typically be retrained for each specific observation, amounting to a large computational overhead. In this work we develop learned post-processing and unrolled iterative methods for varying visibility coverages, proposing training strategies to make these methods agnostic to variations in visibility coverage with minimal to no fine-tuning. Learned post-processing techniques are heavily dependent on the prior information encoded in training data and generalise poorly to other visibility coverages. In contrast, unrolled iterative methods, which include the telescope measurement operator inside the network, achieve good reconstruction quality and computation time, generalising well to other coverages and require little to no fine-tuning. Furthermore, they generalise well to more realistic radio observations and are able to reconstruct images with with a larger dynamic range than the training set.
@@ 1 Introduction
To study radio sources in the Universe at high-resolution there is a need for radio telescopes with a very large aperture that cannot be achieved with a single dish. Instead, radio interferometry forms a telescope using multiple dishes or antennas spaced out over long distances to synthesise an aperture large enough to image the Universe in enough detail to uncover its mysteries. By combining the measurements from two of these antennas, which is referred to as a baseline, a so-called visibility can be measured, corresponding to one Fourier component of the sky. The inversion of this measurement process is an ill-posed problem as the radio interferometric telescope samples the Fourier domain non-uniformly, with a finite number of samples, and are contaminated by measurement noise. By projecting the acquired data to the image domain one recovers what is called a dirty image, a reconstruction of the target convolved with the point-spread function (PSF) of the telescope. Typically, various algorithms are used to reconstruct the image from acquired visibilities to yield a more accurate clean image of the true radio sky (e.g. CLEAN, Högbom, 1974).
@@
Recently, learned approaches have gained in popularity, using training data sets as prior information for reconstructions. One particular approach of these learned methods implements learned regularisation, replacing the conventional regularisers in optimisation methods with learned priors. The most popular and promising method is to use so-called Plug-and-Play (PnP, Venkatakrishnan et al., 2013) denoisers that can be trained on different data domains, foregoing the need for large domain-specific data sets. Since these methods retain the decoupling of the prior and measurement operator they are robust to changes in the measurement operator and measurement noise level. In particular, PnP priors have shown promising results in radio imaging (Terris et al., 2023; Liaudat et al., 2024; Terris et al., 2022; Dabbech et al., 2022; Wilber et al., 2023). Furthermore, recent work has demonstrated how PnP methods can be extended to provide efficient uncertainty quantification (Liaudat et al., 2024). However, these methods require iterative reconstruction and so remain relatively computationally demanding.
@@
Learned unrolled iterative methods typically closely resemble iterative optimisation techniques. These methods unroll a small, fixed number of iterations of the respective optimisation algorithm and replace the proximal operators with learned update steps, typically taking the form of a learned denoiser. Because these methods consist only of a few iterations, with the addition of learned updates, they are much faster than fully iterative approaches. These methods have produced promising results in astronomical imaging (e.g. Mars et al., 2023; Aghabiglou et al., 2024) and strike a balance between speed and quality of reconstruction. Inclusion of the measurement operator in the network architecture allows the network to retain some robustness to changes and uncertainty in the measurement operator (Boink et al., 2019, 2020; Maier et al., 2019), however results are conditioned on the measurement operator used during training.
@@ 2.2 Inverse problem
which provides a trade off between the data fidelity term $\mathcal{L}(\boldsymbol{\Phi}\boldsymbol{x},\boldsymbol{y})$ and the regularisation term $\mathcal{S}(\boldsymbol{x})$, which encodes prior information of the images and is weighted by the regularisation parameter $\lambda$. In radio interferometry, a typical data fidelity term is the squared $\ell_{2}-\text{norm}$, i.e. $\mathcal{L}(\boldsymbol{\Phi}\boldsymbol{x},\boldsymbol{y})=\|\boldsymbol{\Phi}\boldsymbol{x}-\boldsymbol{y}\|_{\ell_{2}}^{2}$, and a widely used regularisation term is an $\ell_{1}$-norm on an overcomplete dictionary of wavelet bases, e.g. consisting of a Dirac basis and Debauchies wavelets bases (Daubechies, 1992), as used in the sparsity averaging reweighted analysis (SARA, Carrillo et al., 2012) algorithm.
The optimisation problem in Equations 7 are typically solved using proximal optimisation techniques, such as the fast iterative shrinkage-thresholding algorithm (FISTA, Beck & Teboulle, 2009), the alternating direction method of multipliers (ADMM, Boyd, 2010), or the primal-dual hybrid gradient method (PDHG, Chambolle et al., 2018) and a variety of these methods have been applied to radio interferometry (e.g. Pratley et al., 2018, 2019a).
Learned reconstruction approaches aim to replace some parts of these traditional optimisation techniques with neural networks in order to incorporate a data-driven prior and/or to speed up the reconstruction process. In learned regularisation approaches neural networks capture the prior information in the variational problem. In particular, PnP methods replace the proximal problem formulation in proximal splitting algorithms with a denoiser (Venkatakrishnan et al., 2013; Ryu et al., 2019).
Learned post-processing methods (e.g. Chen et al., 2017; Jin et al., 2017; Yi & Babyn, 2018) aim to drastically speed up the reconstruction process by training a neural network to post-process an initial reconstruction. Figure 1 shows the typical pipeline of a learned post-processing method in radio interferometry. Because the network is decoupled from the measurement process, these methods can be trained using an image-to-image network making them much faster to train and evaluate.
@@
Learned unrolled iterative methods combine the data-driven priors with model based updates and typically mimic proximal optimisation algorithms by unrolling a small number of iterations of the iterative process and replacing some of the proximal operators with neural networks (Gregor & LeCun, 2010). Figure 2 shows a diagram summarising these methods. While a large portion of these methods closely follow the traditional optimisation methods (e.g. Adler & Öktem, 2018), some methods have been proposed that incorporate the measurement process more directly into the network. For instance, Hauptmann et al. (2020) proposed a multi-scale learned iterative method that includes the measurement information at different resolution scales in the multi-scale U-Net architecture (Ronneberger et al., 2015). Similarly there are methods that see the inclusion of the gradient of a data fidelity term added into this U-Net structure at several resolution scales (Trent, 2020; Mars et al., 2023). The benefit in these methods is the reduced computation and memory overhead as at the lower scales of the network the measurement operator is evaluated at a lower resolution and therefore more efficiently.
@@ 3.2 Learned unrolled iterative method
The second approach we consider is a learned unrolled iterative method that uses the measurement operator and measurements of the observation in the neural network to improve the reconstruction quality. We extend the U-Net architecture by introducing measurement information at each scale of the network. To achieve this we use so-called sub-scale operators $\boldsymbol{\Phi}_{i}:\mathbb{R}^{N}_{i}\rightarrow\mathbb{C}^{M}_{i}$ that map from a reduced image space $\boldsymbol{x}_{i}\in\mathbb{R}^{N_{i}}$ to low-pass filtered visibilities $\boldsymbol{y}_{i}\in\mathbb{C}^{M_{i}}$. These are applied to the down-sampled images at each scale of the network. The Fourier domain of these sub-scale operators is restricted by applying a low-pass filter to the Fourier domain of the full measurement operator, such that the sub-scale operators only contain the measurement information that is relevant for the current scale of the images. The choice of the low-pass filter is discussed in more detail in Mars et al. (2023). These sub-scale operators are used to include the gradient of the data fidelity term $\mathcal{L}(\boldsymbol{\Phi}_{i}\boldsymbol{x}_{i},\boldsymbol{y}_{i})=\frac{1}{2}\|\boldsymbol{\Phi}_{i}\boldsymbol{x}_{i}-\boldsymbol{y}_{i}\|_{\ell_{2}}^{2}$ in the multi-scale U-Net:
@@ 4.2 Reconstruction quality
4.3 Reconstruction quality for increased dynamic range
While the proposed methods are trained on images of simulated galaxies, the resulting networks should also be tested on more realistic radio observations. Therefore, to test the generalisation power of the different training strategies, we evaluate the performance of the models on an out-of-sample image of 30 Doradus. The dynamic range of this image is larger than the simulated images ($\sim 600$) and we therefore also evaluate the reconstructions on a logarithmic scale. While this image has a higher dynamic range than the simulated images, it is not as high as the dynamic range of real radio observations (which can be in the order of $10^{6}$).
@@ 4.3 Reconstruction quality for increased dynamic range
Figure 8: Reconstructions using the trained variants of the U-Net on simulated measurements, using the true uv-coverage drawn from a Gaussian random distribution, of an out-of-sample image of 30 Doradus, displayed on a linear scale (top) and a logarithmic scale (bottom). The strategies that are not fine-tuned on data from the true coverage perform significantly worse in reconstructing this out-of-sample image. Using transfer learning from the single coverage strategy brings the reconstructive SNR closer to fully retraining the model on the true coverage. All U-Net reconstructions struggle to accurately reconstruct the larger dynamic range of this image as indicated by the logarithmically scaled images and the logSNR, with the single coverage strategy producing large artefacts.
When looking at the reconstructions from the same measurements using the GU-Net in Figure 9, we see that the models are able to produce high quality reconstructions with a larger dynamic range. Similar to on the simulated images, the methods not fine-tuned on data from the true coverage performs slightly worse but this difference is easily mitigated by fine-tuning the models using transfer learning, yielding reconstructions that are of the same reconstruction quality as fully retraining the model. The explicit inclusion of the measurement operator in the reconstruction process allows the network to learn the prior information irrespective of the coverage used, which is why the GU-Net is able to generalise much better to out-of-distribution data as well as generalising to different uv-coverages. This generalisablity makes it more suitable for application to real radio observations, since it is able to correct for differences between the training data set and the actual observations.
Figure 9: Reconstructions using the trained variants of the GU-Net on simulated measurements, using the true uv-coverage drawn from a Gaussian random distribution, of an out-of-sample image of 30 Doradus, displayed on a linear scale (top) and a logarithmic scale (bottom). The GU-Net is able to reconstruct the larger dynamic range of the image across the different strategies as indicated by both the high SNR and logSNR of the reconstructions. Training the model on a single coverage and then fine-tuning it using transfer learning provides reconstruction quality close to that of the model that is fully trained on simulated data from the true coverage.
@@ 4.4 Transfer learning to MeerKAT uv-coverage
Finally, to investigate whether these reconstruction networks generalise from the Gaussian visibility coverages to realistic uv-coverages, we use the same models to reconstruct from simulated measurements using a uv-coverage from the MeerKAT telescope with $241920$ visibilities as seen in Figure 10. In order to do this we fine-tune the GU-Net model trained on the single coverage by training it on a small data set created using the MeerKAT coverage. For the creation of the dirty image and the gradient operations inside the GU-Net model, uniform measurement weighting is used. The model is trained for 50 epochs and the creation of the data took $\sim 2$ hours and the fine-tuning of the model took $\sim 5$ hours on an A100 GPU with 40GB VRAM. For this task we only consider the GU-Net architecture as the U-Net models do not generalise well to larger dynamic range images as demonstrated in Section 4.3. The fine-tuned GU-Net is then used to reconstruct an out-of-sample image of 30 Doradus using the MeerKAT coverage. The results are shown in Figure 11.
@@
Figure 11: Reconstructions using a fine-tuned variant of the GU-Net on simulated measurements on an out-of-sample image of 30 Doradus using the MeerKAT uv-coverage as displayed in Figure 10. The ground truth, dirty image and GU-Net reconstruction are displayed on a linear scale (top) and a logarithmic scale (bottom). Using this fine-tuned model we achieve a high quality reconstruction that also reproduces the larger dynamic range of the radio observation, using only minimal amount of re-training.
We see that the transfer learned model is capable of reconstructing images with a larger dynamic range, despite not being trained on any data with a similar dynamic range. Since the number of measurements in the MeerKAT coverage is larger than that of the Gaussian coverage (7.4$\sim$ 7.4$\times$ 7.4 × more measurements), the reconstruction time is longer ($194.6\pm 3.1$ms; 2.3$\sim$ 2.3$\times$ 2.3 × longer) than for the simulated measurements and the training takes longer.
@@ 5 Conclusions
First of all, we conclude that methods that do not include the measurement operator in the reconstruction network, such as learned post-processing methods based on a U-Net model, are not robust to changes in the uv-coverage and are highly reliant on learning features that are part of the training data. While these methods are the most computationally efficient reconstruction approaches, their reconstructive performance is significantly worse than methods that do include the measurement operator, even when fully (re)trained or fine-tuned to the true uv-coverage, due to overfitting. Furthermore, by looking at the performance of this post-processing approach on more realistic radio interferometric observations, we have shown that these methods struggle to capture the larger dynamic range of these images across the different training strategies. This indicates that these methods are limited by the prior information in their training data, which is concerning in the context of real data where the true image is unknown.
@@
Furthermore, we have shown that when applying the GU-Net reconstruction method to measurements outside of the training domain, such as a more realistic radio observation with a larger dynamic range ($\sim 600$), we retain this reconstruction quality, indicative of the approach being significantly less reliant on the prior information in the training data and able to generalise across different types of data.
@@
To extend these method to and test them on real observations, several challenges need to be addressed. For accurate modelling of the measurements, there is need for a version of the radio interferometric measurement operator in a fully automatically differentiable framework that includes wide-field effects, such as w-stacking and w-projection methods. However, using the full measurement operator during training for large uv-coverages would be highly computationally demanding and require a huge amount of memory on the GPU(s) to train for an SKA-sized uv-coverage. In order to address this, we propose to investigate the use of approximate measurement operators in the reconstruction network by using a convolution with the PSF to approximate the measurement operator as a way to reduce the computational overhead of changing the coverage during training, as has been investigated for proximal optimisation methods (Bester et al., 2021). This approximation approach has also been used in Aghabiglou et al. (2024) to train a series of residual networks to reconstruct the image from the measurements. While this approximation does not model wide-field effects when the baselines are non co-planar, this would allow us to train on a distribution of realistic uv-coverages without the need to change the coverage during training, which would be essential in order to train these networks. However, the difference in performance between using the full measurement operator and the approximate operator in learned methods should be investigated further to assess how much information is lost.
While we show that the GU-Net reconstruction models are able to generalise to a larger dynamic range ($\sim 600$) than the training data, we note that this dynamic range is still orders of magnitude lower than the dynamic range of real observations (which can be $\sim 10^{6}$). In order to extend these methods to real observations, more extensive investigation should be performed with regards to creating a representative training data set, and thus prior information, that includes the high dynamic range and bright point sources that are present in real observations, as well as varying noise levels, beam errors, calibration errors and other errors that might be present in the actual data.
An interesting avenue for future work would be to investigate not only the fine tuning on the true uv-coverage, but also the performance of transfer learning on real data. In Mars et al. (2023) we have shown that transfer learning can be used to improve the performance of our reconstruction methods trained on a general image data set, by using a small amount of domain specific data. This approach could bridge the gap between a larger data set of simulated data and a small data set of real data.